# Robust Spacecraft Low-Thrust Trajectory Design: A Chance-Constrained Covariance-Steering Approach

### v1

@@ Abstract
This paper proposes a systematic method for generating practical and robust low-thrust spacecraft trajectories. One contribution is to consider the change in mass of the spacecraft at two levels: a) the propulsive acceleration and b) the intensity of the stochastic disturbances. A covariance variable formulation is considered, which is computationally more efficient than the factorized covariance implementation. The proposed approach is applied to two- (i.e., planar) and three-dimensional heliocentric phases of spacecraft flight from Earth to Mars under the restricted two-body dynamics. The results highlight the importance of keeping track of mass change to generate more realistic, robust solutions for interplanetary space missions to avoid underestimation of mission risks.
@@ I INTRODUCTION
The independence between mean and covariance no longer holds once constraints are present, and thus mean and covariance dynamics become tightly coupled. As a result, standard LQR methods are not sufficient to address the complexity of practical problems. Alternative approaches have sought to integrate LQR with stochastic optimization. One such strategy consists of designing a feedback gain a priori using LQR and then optimizing only the feedforward control component within a stochastic framework [10]. A more flexible formulation treats the feedback gain as an optimization variable, enabling joint design of feedforward and feedback terms to better handle nonlinearities, nonconvex and probabilistic constraints, and covariance propagation. The resulting non-convex problems are typically solved using sequential convex programming (SCP).
@@
The factorized covariance formulation method is proposed in [18, 19, 17] that applies an SCP approach to solve covariance-steering problems. However, factorized covariance formulations are computationally expensive, as each subproblem involves a linear matrix inequality (LMI) whose dimensions grow quadratically with the number of nodes, $N$, leading to a complexity of $\mathcal{O}(N^{2}n_{x}n_{u})$ (with $n_{x}$ and $n_{u}$ denoting the dimensions of state and control vectors, respectively). An alternative covariance variable SCP-based formulation was introduced in [3]. Nevertheless, the proof of lossless convexification underpinning this method was later shown to be incorrect [16]. A subsequent modification by [8] (based on the original work [11]) is also proposed to restore the validity of the formulation under the assumption that a set of tight tolerances are satisfied. Covariance-steering formulations are used for a variety of applications, from powered descent guidance [20] to spacecraft rendezvous and docking maneuvers [25].
Contributions: 1) We investigate the application of the covariance variable SCP-based formulation for solving fuel-optimal low-thrust Earth-to-Mars problems, which explicitly models spacecraft mass dynamics. This formulation a) allows for thrust to be considered as the control input and b) models a mass-varying stochastic force intensity. In previous low-thrust trajectory studies, the mass is not considered as a state. The mass time rate of change is an essential component of the acceleration term that appears in the equations of motion. The advantage of this model is that we can consider realistic engine parameters (such as thrust magnitude and specific impulse) and enforce constraints directly on the maximum thrust that is to be produced by the propulsion system.
@@
2) We also review the steps for implementing a covariance variable formulation, mirroring the steps outlined in Ref. [8], which is computationally more efficient than the factorized covariance implementation. However, we provide particular modifications needed for considering the variation in mass. A detailed analysis of the results is also presented. The paper is structured with Sec. II introducing the general formulation of the stochastic OCP. Section III describes the detailed convexification procedure and the iterative algorithm. The proposed approach is then assessed in Section IV through numerical simulations on a more realistic scenario. Finally, Section V presents concluding remarks with potential directions for future research.
@@
The notation used in this paper is standard. In particular, $\mathbb{E}[.]$ denotes the expectation operator, $\mathbb{P}(\cdot)$ denotes the probability measure, and $\mathbb{S}^{n}_{+}$ is a set of all positive definite matrices of dimension $n$. Also, $x_{t}=x(t)$ and $u_{t}=u(t)$. $\mathbb{R_{+}}$ denotes the set of positive real numbers. $\bm{I}_{n}$ denotes an $n$-dimensional identity matrix.
@@ II-A Continuous-Time Formulation
where $\bar{\bm{x}}_{i},\bar{\bm{x}}_{f}\in\mathbb{R}^{n_{x}}$ are the initial and desired terminal mean states, and ${P}_{i},{P}_{f}\in\mathbb{S}^{n_{x}}_{+}$ are the corresponding covariance matrices. To achieve this, we seek a control policy $\bm{u}_{t}=\pi(\bm{x}_{t},t)$ that simultaneously regulates both the mean state trajectory and the covariance evolution. We employ a piecewise constant affine control of the form
$\bm{u}_{t}=\bm{v}_{t}+{K}_{t}\big(\bm{x}_{t}-\bar{\bm{x}}_{t}\big),\quad t\in[t_{0},t_{f}],$ (3)
where $\bm{v}_{t}\in\mathbb{R}^{n_{u}}$ denotes the nominal control sequence, ${K}_{t}\in\mathbb{R}^{n_{u}\times n_{x}}$ is the feedback gain matrix.
Throughout the time horizon, the state and control trajectories must satisfy various operational and safety requirements. Since the state distribution and consequently the control input, as a state-feedback law, are unbounded under Gaussian assumption, these conditions have to be imposed probabilistically as chance constraints, which can be written as,
@@
where $u_{\max}>0$ is the maximum available control effort and $\beta_{u}$ is its associated confidence level.
While satisfying all constraints, the control objective is also to minimize the control effort. A Lagrange form of the cost functional is written as,
@@ II-B Discrete-Time Deterministic Formulation
where $\Phi(t,\tau)$ is the solution to the matrix ordinary differential equation
@@
The probabilistic control constraint (5) can be reformulated as a conservative deterministic constraint by applying the triangle inequality to the affine control policy as,
$\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\;\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\;\leq\;u_{\max},$ (21)
@@
Similarly, the objective function (7) can be written in a discrete-time deterministic form of
$J_{2}=\sum_{k=0}^{N-1}\Big(\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(p)}\;\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\Big).$ (22)
@@
$\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\leq u_{\max},$ (23d)
@@ III-A Covariance Propagation
To ensure consistency between ${P}_{k}$, ${U}_{k}$, and ${Y}_{k}$, the Schur-complement LMI constraint is imposed
@@ III-B Formulating Chance-Constraint Control
The deterministic equation of the probabilistic control constraint (21) is nonconvex due to square root of the spectral term $\sqrt{\lambda_{\max}({Y}_{k})}$. To convexify, we first introduce another auxiliary variable ${\tau}_{k}\in\mathbb{R}_{+}$ as an upper bound of the maximum standard deviation of the control covariance
@@
where ${\zeta}_{k}\in\mathbb{R}_{+}$ is an additional slack variable to allow for small discrepancies due to linearization. The resulting convex second-order cone inequality constraint becomes
$\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\,\;{\tau}_{k}\;\leq\;u_{\max}.$ (28)
@@ III-C Objective Function
In addition to replacing $\lambda_{\max}({Y}_{k})$ with its upper bound $\bm{\tau}_{k}$ in (22), lossless convexification requires that the gradient with respect to ${Y}_{k}$ be strictly positive [15], $\frac{\partial J}{\partial{Y}_{k}}>0$. This requirement is not satisfied in the current formulation. To guarantee convexity, we augment the cost with a small regularization term involving the trace of ${Y}_{k}$ [8] as,
$J_{3}=\sum_{k=0}^{N-1}\Big(\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(p)}\,\;{\tau}_{k}+\varepsilon_{Y}\,\mathrm{tr}({Y}_{k})\Big).$ (29)
@@
$J_{\mathrm{pen}}=\sum_{k=0}^{N-1}\Big({\zeta}_{k}+\tfrac{w}{2}{\zeta}_{k}^{2}+\sqrt{w}\,\|{\zeta}\|_{1}\big),$ (30)
@@
Problem 3 (Convex SDP Covariance-Steering Problem). Let $\bm{z}_{3}={\{\bar{\bm{x}},\bm{v}_{k},{P}_{k},{U}_{k},{Y}_{k},{\tau}_{k},{\zeta}_{k}\}}$, the convex SDP problem, which is solved at each SCP iteration, is written as,
@@
$\|\bm{v}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\,\;\bm{\tau}_{k}\;\leq\;u_{\max},$ (31e)
$\lambda_{\max}({Y}_{k})-\bm{\hat{\tau}}_{k}^{2}-2\bm{\hat{\tau}}_{k}(\bm{\tau}_{k}-\bm{\hat{\tau}}_{k})\;\leq\;\bm{\zeta}_{k},$ (31f)
@@
    Initialization: choose a nominal feasible trajectory $\{\bar{\bm{x}}_{k},\bm{v}_{k},\bm{\hat{\tau}}_{k}\}$ and select penalty weights $\varepsilon_{Y},w>0$.
@@
    Solve Convex SDP: solve problem 3 to obtain $\bm{z}=\{\bar{\bm{x}}_{k}^{*},\bm{v}_{k}^{*},{P}_{k}^{*},{U}_{k}^{*},{Y}_{k}^{*},\bm{\tau}_{k}^{*},\bm{\zeta}_{k}^{*}\}$.
@@
    Iterate until the solution converges.
@@ IV Numerical Simulations
IV-A Planar low-thrust Earth-to-Mars problem
@@
where $\bm{x}_{t}=[\,\bm{r}_{t}^{\top},\,\bm{v}_{t}^{\top},\,m_{t}\,]^{\top}\in\mathbb{R}^{5}$ is the state vector, with $\bm{r}_{t}\in\mathbb{R}^{2}$ the position, $\bm{v}_{t}\in\mathbb{R}^{2}$ the velocity, and $m_{t}\in\mathbb{R}$ the spacecraft mass. The control input, $\bm{u}_{t}\in\mathbb{R}^{2}$, represents the thrust vector. The parameter $\mu$ denotes the Sun’s gravitational parameter, and $I_{\text{sp}}$ and $g_{0}$ are the specific impulse and Sea-level gravity values, respectively. We consider the inclusion of mass because its change significantly influences both the optimal trajectory and the stochastic behavior of the system, unlike many simplified models [19, 3]. In addition, we can consider the thrust magnitude as the control input, which is more practical as opposed to considering maximum acceleration [12]. The corresponding SDEs become
@@
where $d\bm{w}_{t}\in\mathbb{R}^{2}$ is a two-dimensional Wiener process with independent increments and $\gamma$ is the force disturbance intensity. The diffusion term scales $\gamma$ by the inverse of the instantaneous spacecraft mass. Thus, as $m_{t}$ decreases, the same level of force uncertainty induces a larger stochastic acceleration disturbance. The continuous-time dynamics are linearized according to Eqs. (10-11) and subsequently discretized by Eqs. (13-17) over $N=40$ time intervals. The initial reference trajectory, required for linearization, is obtained by solving a minimum-fuel deterministic version of the problem using the CasADi solver [1] following the work in [12]. The boundary conditions, along with the relevant physical and mission parameters, are summarized in Table I.
@@
Although the chosen disturbance intensity (i.e., $\gamma/m(t_{0})$) is comparable to the magnitude reported in the literature [19, 3], the effective disturbance level exceeds those values by the end of the maneuver due to the decreasing spacecraft mass. Note that $m(t_{0})$ is assumed to be deterministic at departure and a large terminal mass covariance is permitted, reflecting the fact that the control objective does not involve regulating the final mass uncertainty.
To avoid numerical conditioning issues during optimization, all problem data are properly scaled. In particular, the covariance dynamics and related constraints are scaled following the approach proposed in [8], with a scaling factor of $d$. The SCP algorithm terminates when the relative difference between consecutive state trajectories satisfies
@@
where $(i)$ denotes the SCP iteration and the linearization slack variable satisfies $\zeta\leq\varepsilon_{\zeta}=10^{-6}.$ Penalty coefficient $w$ is updated at each iteration, $i$, as,
@@
Problem 3 is solved and the convergence is achieved in 12 iterations (6.01 s) on an Intel® Core™ i7-13620H @ 2.40 GHz laptop computer. Figure 1 depicts the evolution of the position covariance along with the mean trajectory. In the open-loop case, the covariance grows rapidly over time, whereas in the closed-loop case, the $95\%$-confidence ellipses remain bounded due to the action of the feedback controller.
(a) Open-loop position covariance.
(b) Closed-loop position covariance.
Figure 1: Position covariances.
@@
A Monte Carlo simulation (with 1000 samples) was used to validate the optimized trajectory. Figure 2 compares the resulting dispersion with the nominal solution in which both position and velocity covariance ellipses in Figs. 2(a) and 2(b) are enlarged tenfold for improved visual interpretation. The close alignment of position and velocity samples within their predicted uncertainty ellipses validates the precision of the covariance propagation framework.
@@
Figure 3 depicts the control profiles and its components. Figure 3(a) shows the closed-loop nominal control ($\bm{v}$) and Monte Carlo simulation (1000 samples) profiles. While the feedforward control consists of three main thrust arcs (similar to the deterministic solution for the same problem [22]), the magnitude of the thrust shows noticeable deviations from a clean bang-bang profile. This is a direct consequence of trading off robustness for pure fuel optimality, which requires the mean thrust to create some margins for the feedback term to control the covariances, which has occurred mainly during the last two thrust arcs. Figure 3(b) shows the components of the nominal control with feedback gain uncertainty ellipses.
(a) Closed-loop position covariances.
(b) Closed-loop velocity covariances.
Figure 2: Position covariances (scaled 10x).
@@
(a) Magnitude of the Monte Carlo control samples (gray) and closed-loop nominal control (black).
@@
(b) Control with feedback gain uncertainty ellipses.
Figure 3: Control mean and covariance evolution.
@@
Figure 4 shows the time histories of the mass and its standard deviation along the trajectory. The mass decreases by approximately $25\%$ over the course of the mission, reaching a final value of approximately $m_{f}=3775$ kg upon arrival at Mars. The velocity standard deviation, shown in Figure 5, exhibits a significant growth over the mission horizon when the mass uncertainty is included, with peak standard deviations increasing by up to $16.79\%$. A similar trend is expected in the evolution of the position uncertainty and, as Figure 6 shows the trace of the position covariances, mass inclusion results in $33.27\%$ higher dispersion at mid-course phase. These differences underscore the indirect amplification mechanism by which uncertainty in mass, originating from stochastic control input norms, propagates through the dynamics and inflates the velocity variance due to the $1/m$ coupling in the translational acceleration equation. Neglecting this effect leads to an underestimation of the control authority required to robustly satisfy the terminal state covariance constraints, especially in long-duration, low-thrust missions.
Figure 4: Mass and its standard deviation vs. time.
@@
Figure 5: Velocity standard deviation along first principle component with (black) and without (red) mass uncertainty.
@@
Figure 6: Trace of the position covariances with (black) and without (red) the inclusion of mass uncertainty.
Figure 7: Control w/ (black) and w/o (red) mass uncertainty.

### v2

@@ Abstract
This paper proposes a systematic method for generating practical and robust low-thrust spacecraft trajectories. One contribution is to consider the change in mass of the spacecraft at two levels: a) the propulsive acceleration and b) the intensity of the stochastic disturbances. A covariance variable formulation is considered, which is computationally more efficient than the factorized covariance implementation. The proposed approach is applied to two- (i.e., planar) and three-dimensional heliocentric phases of spacecraft flight from Earth to Mars under the restricted two-body dynamics. The results highlight the importance of keeping track of mass change to generate more realistic, robust trajectories for interplanetary space missions to avoid underestimation of mission risks.
@@ I INTRODUCTION
The independence between mean and covariance no longer holds once constraints are present, and thus mean and covariance dynamics become tightly coupled. As a result, standard LQR methods are not sufficient to address the complexity of practical problems. Alternative approaches have sought to integrate LQR with stochastic optimization. One such strategy consists of designing a feedback gain a priori using LQR and then optimizing only the feedforward control component within a stochastic framework [11]. A more flexible formulation treats the feedback gain as an optimization variable, enabling joint design of feedforward and feedback terms to better handle nonlinearities, nonconvex and probabilistic constraints, and covariance propagation. The resulting non-convex problems are typically solved using sequential convex programming (SCP) [12].
@@
The factorized covariance formulation method is proposed in [13, 14, 15] that applies an SCP approach to solve covariance-steering problems. However, factorized covariance formulations are computationally expensive, as each subproblem involves a linear matrix inequality (LMI) whose dimensions grow quadratically with the number of nodes, $N$, leading to a complexity of $\mathcal{O}(N^{2}n_{x}n_{u})$ (with $n_{x}$ and $n_{u}$ denoting the dimensions of state and control vectors, respectively). An alternative covariance variable SCP-based formulation was introduced in [16]. Nevertheless, the proof of lossless convexification underpinning this method was later shown to be incorrect [17]. A subsequent modification by [18] (based on the original work [19]) is also proposed to restore the validity of the formulation under the assumption that a set of tight tolerances are satisfied. Covariance-steering formulations are used for various applications, including path planning [20], powered descent guidance [21], aerocapture [22] and spacecraft rendezvous and docking maneuvers [23].
@@
Contributions: 1) We investigate the application of the covariance variable SCP-based formulation for solving fuel-optimal low-thrust Earth-to-Mars problems, which explicitly models spacecraft mass dynamics. This formulation a) allows for thrust to be considered as the control input, b) models a mass-varying stochastic force intensity, and c) allows for enforcing a standard deviation in spacecraft final mass, which ties it nicely with some of the recent research on desensitization of spacecraft trajectories [24]. In previous applications of covariance steering for low-thrust trajectory design, the mass is not considered as a state. However, the instantaneous mass is an essential component of the propulsive acceleration term that appears in the equations of motion. The advantage of this model is that we can consider realistic engine parameters (such as thrust magnitude and specific impulse) and enforce constraints directly on the maximum thrust that is to be produced by the propulsion system. 2) We also review the steps for implementing a covariance variable formulation, mirroring the steps outlined in Ref. [18], which is computationally more efficient than the factorized covariance implementation. However, we provide particular modifications needed for considering the variation in mass. A detailed analysis of the results is also presented. The paper is structured with Sec. II introducing the general formulation of the stochastic OCP. Section III describes the detailed convexification procedure and the iterative algorithm. The proposed approach is then assessed in Section IV through numerical simulations on more realistic scenarios. Finally, Section V presents concluding remarks with potential directions for future research.
@@
The notation used in this paper is standard. In particular, $\mathbb{E}[.]$ denotes the expectation operator, $\mathbb{P}(\cdot)$ denotes the probability measure, and $\mathbb{S}^{n}_{+}$ is a set of all positive definite matrices of dimension $n$. Also, $x_{t}=x(t)$ and $u_{t}=u(t)$. $\mathbb{R^{+}}$ denotes the set of positive real numbers. $\bm{I}_{n}$ denotes an $n$-dimensional identity matrix.
@@ II-A Continuous-Time Formulation
where $\bar{\bm{x}}_{i},\bar{\bm{x}}_{f}\in\mathbb{R}^{n_{x}}$ are the initial and desired (final) mean states, and ${P}_{i},{P}_{f}\in\mathbb{S}^{n_{x}}_{+}$ are the corresponding covariance matrices. To achieve this, we seek a control policy $\bm{u}_{t}=\pi(\bm{x}_{t},t)$ that simultaneously regulates both the mean state trajectory and its covariance evolution. We employ a piecewise constant affine control of the form
$\bm{u}_{t}=\bm{F}_{t}+{K}_{t}\big(\bm{x}_{t}-\bar{\bm{x}}_{t}\big),\quad t\in[t_{0},t_{f}],$ (3)
where $\bm{F}_{t}\in\mathbb{R}^{n_{u}}$ denotes the nominal (feed-forward) control sequence, ${K}_{t}\in\mathbb{R}^{n_{u}\times n_{x}}$ is the feedback gain matrix.
Throughout the time horizon, the state and control trajectories must satisfy various operational and safety requirements. Since the state distribution and consequently the control input, as a state-feedback law, are unbounded under Gaussian assumptions, these conditions have to be imposed probabilistically as chance constraints, which can be written as,
@@
where $u_{\max}>0$ is the maximum available control effort and $\beta_{u}$ is its associated confidence level. Please refer to Ref. [25] for incorporating hard constraints into covariance-steering problems.
While satisfying all constraints, the control objective is also to minimize the cumulative control effort. A Lagrange form of the cost functional is written as,
@@ II-B Discrete-Time Deterministic Formulation
where $\Phi(t,\tau)$ is the solution to the matrix ordinary differential equation [26] as,
@@
The probabilistic control constraint (5) can be reformulated as a conservative deterministic constraint by applying the triangle inequality to the affine control input as,
$\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\;\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\;\leq\;u_{\max},$ (21)
@@
Similarly, the objective function (7) can be written in a discrete-time deterministic form as,
$J_{2}=\sum_{k=0}^{N-1}\Big(\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(p)}\;\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\Big).$ (22)
@@
$\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\sqrt{\lambda_{\max}({K}_{k}{P}_{k}{K}_{k}^{\top})}\leq u_{\max},$ (23d)
@@ III-A Covariance Propagation
To ensure consistency between ${P}_{k}$, ${U}_{k}$, and ${Y}_{k}$, the Schur-complement LMI constraint is imposed as,
@@ III-B Formulating Chance-Constraint Control
The deterministic equation of the probabilistic control constraint (21) is nonconvex due to the square root of the spectral term $\sqrt{\lambda_{\max}({Y}_{k})}$. To convexify, we first introduce another auxiliary variable ${\tau}_{k}\in\mathbb{R}^{+}$ as an upper bound of the maximum standard deviation of the control covariance
@@
where ${\zeta}_{k}\in\mathbb{R}^{+}$ is an additional slack variable to allow for small discrepancies due to linearization. The resulting convex second-order cone inequality constraint becomes
$\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\,\;{\tau}_{k}\;\leq\;u_{\max}.$ (28)
@@ III-C Objective Function
In addition to replacing $\lambda_{\max}({Y}_{k})$ with its upper bound $\tau_{k}$ in (22), lossless convexification requires that the gradient with respect to ${Y}_{k}$ be strictly positive [27], i.e., $\frac{\partial J}{\partial{Y}_{k}}>0$. This requirement is not satisfied in the current formulation. To guarantee convexity, we augment the cost with a small regularization term involving the trace of ${Y}_{k}$ [18] as,
$J_{3}=\sum_{k=0}^{N-1}\Big(\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(p)}\,\;{\tau}_{k}+\varepsilon_{Y}\,\mathrm{tr}({Y}_{k})\Big).$ (29)
@@
$J_{\mathrm{pen}}=\sum_{k=0}^{N-1}\Big({\zeta}_{k}+\tfrac{w}{2}{\zeta}_{k}^{2}+\sqrt{w}\,\|{\zeta_{k}}\|_{1}\big),$ (30)
@@
Problem 3 (Convex SDP Covariance-Steering Problem). Let $\bm{z}_{3}={\{\bar{\bm{x}}_{k},\bm{F}_{k},{P}_{k},{U}_{k},{Y}_{k},{\tau}_{k},{\zeta}_{k}\}}$, the convex SDP problem (solved within each SCP iteration) is written as,
@@
$\|\bm{F}_{k}\|+\sqrt{Q_{\chi^{2}_{n_{u}}}(\beta_{u})}\,\;\tau_{k}\;\leq\;u_{\max},$ (31e)
$\lambda_{\max}({Y}_{k})-\hat{\tau}_{k}^{2}-2\hat{\tau}_{k}(\tau_{k}-\hat{\tau}_{k})\;\leq\;\zeta_{k},$ (31f)
@@
    Initialization: choose a nominal feasible trajectory $\{\bar{\bm{x}}_{k},\bm{F}_{k},\hat{\tau}_{k}\}$ and select penalty weights $\varepsilon_{Y},w>0$.
@@
    Solve Convex SDP: solve problem 3 to obtain $\bm{z}=\{\bar{\bm{x}}_{k}^{*},\bm{F}_{k}^{*},{P}_{k}^{*},{U}_{k}^{*},{Y}_{k}^{*},\tau_{k}^{*},\zeta_{k}^{*}\}$.
@@
    Iterate until the solution converges based on $\varepsilon_{x}$ and $\varepsilon_{\zeta}$.
@@ IV Numerical Simulations
IV-A Planar (2D) low-thrust Earth-to-Mars problem
@@
$\bm{f}(\bm{x}_{t},\bm{u}_{t},t)=\begin{bmatrix}\bm{v}_{t}\\[6.0pt] -\mu\dfrac{\bm{r}_{t}}{\|\bm{r}_{t}\|^{3}}+\dfrac{\bm{u}_{t}}{m_{t}}\\[10.0pt] -\dfrac{\|\bm{u}_{t}\|}{I_{\text{sp}}g_{0}}\end{bmatrix},~\text{with}~\bm{u}_{t}=[u_{1},u_{2}]^{\top}$ (32)
where $\bm{x}_{t}=[\,\bm{r}_{t}^{\top},\,\bm{v}_{t}^{\top},\,m_{t}\,]^{\top}\in\mathbb{R}^{5}$ is the state vector, with $\bm{r}_{t}\in\mathbb{R}^{2}$ the position, $\bm{v}_{t}\in\mathbb{R}^{2}$ the velocity, and $m_{t}\in\mathbb{R^{+}}$ the spacecraft mass. The control input, $\bm{u}_{t}\in\mathbb{R}^{2}$, represents the thrust vector. The parameter $\mu$ denotes the Sun’s gravitational parameter, and $I_{\text{sp}}$ and $g_{0}$ are the specific impulse and sea-level gravity values, respectively. We consider the inclusion of mass because its change significantly influences both the optimal trajectory and the stochastic behavior of the system, unlike some previous studies [14, 16, 18]. In addition, we can consider the thrust magnitude as the control input, which is more practical as opposed to considering maximum acceleration. We note, however, that formulating and generating spacecraft trajectories based on propulsive acceleration is a reasonable assumption, in particular, for electric low-thrust propulsion systems and over short time horizons [29] or for preliminary studies [30]. The corresponding SDEs become
@@
where $d\bm{w}_{t}\in\mathbb{R}^{2}$ is a two-dimensional Wiener process with independent increments and $\gamma$ is the force disturbance intensity. The diffusion term scales $\gamma$ by the inverse of the instantaneous spacecraft mass. As $m_{t}$ decreases, the same level of force uncertainty induces a larger stochastic acceleration disturbance. The continuous-time dynamics are linearized according to Eqs. (10)-(11) and subsequently discretized by Eqs. (13)-(17) over $N=40$ time intervals. The initial reference trajectory, required for linearization, is obtained by solving a minimum-fuel deterministic version of the problem using the CasADi solver [31] following the work in [32]. The boundary conditions, along with the relevant physical and mission parameters, are summarized in Table I.
@@
Although the chosen disturbance intensity (i.e., $\gamma/m(t_{0})$) is comparable to the magnitude reported in the literature [14, 16], the effective stochastic disturbance level exceeds those values by the end of the maneuver due to the decreasing spacecraft mass. Note that $m(t_{0})$ is assumed to be deterministic at departure, and a large terminal mass variance is permitted, reflecting the fact that the control objective does not involve regulating the final mass uncertainty (i.e., variance).
To avoid numerical conditioning issues during optimization, all problem data are properly scaled. In particular, the covariance dynamics and related constraints are scaled with a scaling factor of $d$ [18]. The SCP algorithm terminates when the relative difference between consecutive state trajectories satisfies the following constraint as,
@@
where $(i)$ denotes the SCP iteration and the linearization slack variable satisfies $\zeta\leq\varepsilon_{\zeta}=10^{-6}$. The penalty coefficient $w$ is updated at the $i$-th iteration as,
@@
Problem 3 is solved, and the convergence is achieved in 12 iterations (6.01 s) on an Intel® Core™ i7-13620H @ 2.40 GHz laptop computer. Figure 1 depicts the evolution of the position covariance along with the mean trajectory. In the open-loop case, the covariance grows rapidly over time, whereas in the closed-loop case, the $95\%$-confidence ellipses remain bounded due to the action of the feedback controller.
(a) Open-loop position covariance ellipses.
(b) Closed-loop position covariance ellipses.
Figure 1: 2D case: open-loop and closed-loop trajectories along with their position covariance ellipses.
@@
A Monte Carlo simulation (with 1000 samples) was used to validate the optimized trajectory. Figure 2 compares the resulting dispersion with the nominal solution, in which both position and velocity covariance ellipses in Figs. 2(a) and 2(b) are enlarged tenfold for improved visual interpretation. The close alignment of position and velocity samples within their predicted uncertainty ellipses validates the precision of the covariance propagation framework.
(a) Closed-loop position covariance ellipses.
(b) Closed-loop velocity covariance ellipses.
Figure 2: 2D case: position and velocity covariance ellipses (scaled 10x) of the closed-loop solution.
@@
Figure 3 depicts the control profiles and their components. Figure 3(a) shows the feed-forward control ($\|\bm{F}\|$) and Monte Carlo simulation (1000 samples) profiles. While the feedforward control consists of three main thrust arcs (similar to the deterministic solution for the same problem [28]), the magnitude of the thrust shows noticeable deviations from a clean bang-bang profile. This is a direct consequence of trading off pure fuel optimality for trajectory robustness, which requires the mean thrust to create some margins for the feedback term to control the covariances, which has occurred mainly during the last two thrust arcs. Figure 3(b) shows the components of the nominal control with feedback gain uncertainty ellipses.
@@
(a) Magnitude of the feedforward control (black) and Monte Carlo control samples (gray).
@@
(b) Feedforward control, denoted as ‘+’ with feedback gain uncertainty ellipses.
Figure 3: 2D case: control mean and covariance ellipses.
@@
Figure 4 shows the time histories of the mass and its standard deviation along the trajectory (which is calculated from the covariance solution matrix at the final point). The mass decreases over the course of the mission, reaching a final value of approximately $m_{f}=3774.59$ kg upon arrival at Mars. The velocity standard deviation, shown in Figure 5, exhibits a significant growth over the mission when the mass uncertainty is included, with peak standard deviations increasing by up to $16.79\%$. A similar trend is expected in the evolution of the position uncertainty, and as Figure 6 shows the trace of the position covariances, the inclusion of mass results in $33.27\%$ higher dispersion at the mid-course phase. These differences underscore the indirect amplification mechanism by which uncertainty in mass, originating from stochastic control input norms, propagates through the dynamics and inflates the velocity variance due to the $1/m$ coupling in the translational acceleration equation. Neglecting this effect leads to an underestimation of the control authority required to robustly satisfy the terminal state covariance constraints, especially in long-duration, low-thrust missions.
Figure 4: 2D case: mass and its standard deviation vs. time.
@@
Figure 5: 2D case: velocity standard deviation along the first principal component with (black) and without (red) mass variations.
@@
Figure 6: 2D case: trace of the position covariances with (black) and without (red) the inclusion of mass variations.
Figure 7: 2D case: control w/ (black) and w/o (red) mass uncertainty.